OpenStudy (anonymous):

What is the value of cos 17 cos 81 + sin 43 cos 69 + 1/2 cos 142?

OpenStudy (anonymous):

Could you show us what you've done so far please?

OpenStudy (anonymous):

Yes, it is the original question :-( Maybe you can calculate cos 17 or cos 81 or sin 43 or cos 69 or cos 142?

OpenStudy (anonymous):

IF not, you can just use a calculator and find the value to the nearest 2 decimal places or something.

OpenStudy (anonymous):

But I'm guessing RolyPoly is right about needing the exact value.

OpenStudy (anonymous):

Yes, but without calculator

OpenStudy (anonymous):

Well first I can see that sin43 can be rewritten as cos(90-43)

OpenStudy (anonymous):

Then cos(90-43)=cos(47) Which can be rewritten again as cos(30+17)

OpenStudy (anonymous):

Then you use angle results.

OpenStudy (anonymous):

ANd then cos81 can be rewritten as cos(150-69) and you can use angle results on that. And you know that: \[\cos150=-\frac{\sqrt{3}}{2}\] and: \[\sin150=\frac{1}{2}\]

OpenStudy (anonymous):

@RolyPoly I agree that would work as well, but you would just form another different angle which is 21, so it would be a bother to find that as well.

OpenStudy (anonymous):

Then?

OpenStudy (anonymous):

So your equation becomes this: \[\cos17[\cos(150-69) + \cos(30+17)\cos69 + \frac{1}{2}\cos142=?\] COuld you expand the double angle results please.

OpenStudy (anonymous):

Everything will cancel, when you expand everything out the way I planned it for you.

OpenStudy (anonymous):

And then you would be left with: \[\frac{1}{2}\cos142\] What I need you to do is show me that you understand YOUR OWN question.

OpenStudy (anonymous):

You don't just let me do the talking. You need to really do it yourself. I think I've done 3/4 of it. You should be capable of going on. I will be here to guide you from now on. So you should show me what you've done from here on then.

OpenStudy (anonymous):

My answer is still at \[\frac{ 1 }{ 2 } (\sin 69 \cos 17 + \cos 69 \sin 17) + \frac{ 1 }{ 2 } \cos 142\] \[\frac{ 1 }{ 2 } (\sin 86 + \cos 142)\] Then what should i do?

OpenStudy (anonymous):

@RolyPoly and @fiedar I will scan the working out for you UP to this stage and you can see everything cancels out.

OpenStudy (anonymous):

Ok, thank you. I will wait for it

OpenStudy (anonymous):

It looks like you've gone ahead with it. Yeah you've gotten it right. Just a mistake at your second last line. It's meant to be a minus sign inside those brackets/paranthesis.

OpenStudy (anonymous):

And then what do you think we should do?

OpenStudy (anonymous):

Oh, I see. it must be negative. Then?

OpenStudy (anonymous):

Nah it's just a minor mistake. You're fine with the last line. Just the second last line.

OpenStudy (anonymous):

So when you have sin86 it can be written as sin(60+26). ANd then I can see something happening with cos142. It can be written as cos(90+52). ANd the angle 52 is twice that of 26. So it would give you an idea of what to do now.

OpenStudy (anonymous):

YOu would have to expand cos142 twice due to the double angle of 52.

OpenStudy (anonymous):

Oh, OK. Thank you for your help :D

OpenStudy (anonymous):

Now worries man, I will put up my scanned working out if you need to look on something.

OpenStudy (anonymous):

Wait something's wrong.

OpenStudy (anonymous):

Sorry you got the second line and last line wrong.

OpenStudy (anonymous):

I was right about the minus sing but the last line should be 1/2(sin52 +cos142)

OpenStudy (anonymous):

minus sign*

OpenStudy (anonymous):

so you wouldn't need to do much with the double angle results.

OpenStudy (anonymous):

Yes. Thank you once more

OpenStudy (anonymous):

And you're right @RolyPoly Both methods are correct. Well Done mate.